# Towards Anomaly-Aware Pre-Training and Fine-Tuning for Graph Anomaly Detection

### v2

@@ 5 Conclusion
This paper tackles label scarcity and homophily disparity in GAD by introducing APF. APF leverages Rayleigh Quotient-guided subgraph sampling and dual spectral filters to capture both semantic and anomaly-sensitive signals without supervision. During fine-tuning, a node- and dimension-adaptive fusion mechanism, together with anomaly-aware regularization, enhances the model’s ability to distinguish abnormal nodes under homophily disparity. Both theoretical analysis and extensive experiments on 10 benchmark datasets validate the effectiveness of our approach.
Acknowledgments and Disclosure of Funding
We would like to express our sincere gratitude to the reviewers and chairs for their invaluable time and effort dedicated to the evaluation of our manuscript.
Appendix A Detailed Preliminaries
@@
In a general scenario, we are given an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\boldsymbol{X})$, where $\mathcal{V}=\{v_{1},\cdots,v_{n}\}$ is the set of $n$ nodes, =$\mathcal{E}=\{e_{ij}\}$ is the set of edges, and $e_{ij}=(v_{i},v_{j})$ represents an edge between nodes $v_{i}$ and $v_{j}$. We define $\boldsymbol{A}$ as the corresponding adjacency matrix and $\boldsymbol{D}$ as the diagonal degree matrix with $\boldsymbol{D}_{ii}=d_{i}=\sum_{j}\boldsymbol{A}_{ij}$. The neighbor set $\mathcal{N}_{i}$ of each node $v_{i}$ is given by $\mathcal{N}_{i}=\{v_{j}:e_{ij}\in\mathcal{E}\}$. For each node $v_{i}$, it has a $d$-dimensional feature vector $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$, and collectively the features of all nodes are denoted as $\boldsymbol{X}=(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n})^{\top}\in\mathbb{R}^{n\times d}$.
@@
GAD is formulated as a binary classification problem where anomalies are regarded as positive with label 1, while normal nodes are negative with label 0. Given $\mathcal{V}^{L}=\mathcal{V}_{a}\cup\mathcal{V}_{a}$, where $\mathcal{V}_{a}$ consists of labeled abnormal nodes and $\mathcal{V}_{n}$ comprises labeled normal nodes, along with their corresponding labels $\boldsymbol{y}^{L}$, the goal is to identify the anomalous status $\boldsymbol{\hat{y}}^{U}$ for the unlabeled nodes $\mathcal{V}^{U}=\mathcal{V}\setminus\mathcal{V}^{L}$. Usually, obtaining authentic labels is often costly, we assume that label information is only available for a small subset of nodes (i.e., $|\mathcal{V}^{L}|\ll|\mathcal{V}|$). Furthermore, there are usually significantly fewer abnormal nodes than normal nodes (i.e., $|\mathcal{V}_{a}|\ll|\mathcal{V}_{n}|$).
@@
Given the adjacency matrix $\boldsymbol{A}$, the graph Laplacian matrix is defined as $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{A}$, which is symmetric and positive semi-definite. Their symmetric normalized versions are noted as $\tilde{\boldsymbol{A}}=\boldsymbol{D}^{-1/2}\boldsymbol{A}\boldsymbol{D}^{-1/2}$ and $\tilde{\boldsymbol{L}}=\boldsymbol{I}-\boldsymbol{D}^{-1/2}\boldsymbol{A}\boldsymbol{D}^{-1/2}$. Its eigendecomposition is given by $\tilde{\boldsymbol{L}}=\boldsymbol{U}\boldsymbol{\boldsymbol{\Lambda}}\boldsymbol{U}^{\top}$, where the columns of $\boldsymbol{U}\in\mathbb{R}^{n\times n}$ are orthonormal eigenvectors (graph Fourier basis), and $\boldsymbol{\boldsymbol{\Lambda}}=\operatorname{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ contains the eigenvalues (frequencies), arranged such that $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$. Given the graph features $\boldsymbol{X}$ and a filter function $g(\cdot)$, the corresponding filtered features is thus defined as $\boldsymbol{\hat{X}}=\boldsymbol{U}g(\boldsymbol{\boldsymbol{\Lambda}})\boldsymbol{U}^{\top}\boldsymbol{X}$.
@@
Typically, $\tilde{\boldsymbol{A}}$ acts as a low-pass filter with $g(\lambda)=1-\lambda$, while $-\tilde{\boldsymbol{A}}$ and $\tilde{\boldsymbol{L}}$ serve as high-pass filters, with $g(\lambda)=\lambda-1$ and $g(\lambda)=\lambda$, respectively. In practice, a self-loop is often added to each node in the graph (i.e., $\boldsymbol{A}=\boldsymbol{A}+\boldsymbol{I}$) to alleviate numerical instabilities and improve performance [34].
Appendix B Related Works
B.1 Graph Anomaly Detection
@@
GNN-based approaches have emerged as a promising paradigm for GAD, due to their strong ability to capture both complex structural and node attribute patterns in graph data. A comprehensive and up-to-date survey of deep GAD methods is provided in [53], while BOND [39] and GADBench [59] establish performance benchmarks for unsupervised and semi-/supervised GAD approaches, respectively.
@@
Unsupervised GAD approaches do not rely on labeled data for training and instead use unsupervised learning techniques to identify anomaly patterns in graph data. For instance, DOMINANT [16] employs a graph autoencoder to reconstruct both attributes and structure using GNNs. CoLA [41] explores the consistency between anomalies and their neighbors across different contrastive views to assess node irregularity. VGOD [31] combines variance-based and attribute reconstruction models to detect anomalies in a balanced manner. TAM [52] introduces local affinity as an anomaly measure, aiming to learn tailored node representations for GAD by maximizing the local affinity between nodes and their neighbors.
@@
Semi-/supervised GAD approaches assume that labels for some normal and abnormal nodes are available for training. They aim to assign labels by learning a decision boundary between normal and abnormal nodes. For example, BWGNN [60] uses a Beta graph wavelet to learn band-pass filters that capture anomaly signals. AMNet [6] employs a restricted Bernstein polynomial parameterization to approximate filters in multi-frequency groups. CARE-GNN [19], PCGNN [40], and GHRN [21] adaptively prune inter-class edges based on neighbor distributions or the graph spectrum. PMP [74] introduces a partitioned message-passing mechanism to handle homophilic and heterophilic neighbors independently. To address settings with limited labeled data, CGenGA [46] proposes a diffusion-based graph generation method to synthesize additional training nodes, while ConsisGAD [8] incorporates learnable data augmentation to utilize the abundance of unlabeled data for consistency training. Furthermore, GGAD [54] introduces a novel semi-supervised framework using only labeled normal nodes.
@@
Additionally, several works have explored GAD within the pre-training and fine-tuning paradigm. DCI [67] decouples representation learning and classification through a cluster-enhanced self-supervised learning task. [12] evaluates the performance of DGI [63] and GraphMAE [29] for GAD, demonstrating the potential of leveraging graph pre-training to enhance GAD with limited supervision. However, most existing methods pre-train a uniform global low-pass filter (e.g., GCN [34]) and then fine-tune a classifier on frozen node representations. The heterophily of abnormal nodes in GAD presents a significant challenge for directly applying these methods. To address this, we propose a pre-training and fine-tuning framework tailored for GAD.
B.2 Graph Pre-Training
@@
Graph pre-training has emerged as a promising paradigm for label-efficient learning [32]. These methods first learn universal knowledge from unlabeled data using self-supervised objectives, which are then transferred to tackle specific downstream tasks. Existing pre-training approaches can be broadly categorized into two groups: contrastive and non-contrastive approaches.
@@
Contrastive approaches typically follow the principle of mutual information maximization [27], where the objective functions contrast positive pairs against negative ones. For instance, DGI [63] and DCI [67] focus on representation learning by maximizing the mutual information between node-level representations and a global summary representation. GRACE [72], GCA [73], and NS4GC [42] learn node representations by pulling together the representations of the same node (positive pairs) across two augmented views, while pushing apart the representations of different nodes (negative pairs) across both views.
@@
Non-contrastive approaches, on the other hand, eliminate the need for negative samples. For example, CCA-SSG [70] and G-BT [3] aim to learn augmentation-invariant information while introducing feature decorrelation to capture orthogonal features. BGRL [61] and BLNN [44] employ asymmetric architectures that learn node representations by predicting alternative augmentations of the input graph and maximizing the similarity between the predictions and their corresponding targets. GraphMAE [29] focuses on feature reconstruction using a masking strategy and scaled cosine error. Additionally, SSGE [43] minimizes the distance between the distribution of learned representations and the isotropic Gaussian distribution to promote the uniformity of node representations.
@@
However, the methods discussed above rely on low-pass GNN encoders that inherently smooth neighbor representations, leading to unsatisfactory performance on heterophilic abnormal nodes. Although a recent work, PolyGCL [7], employs both low- and high-pass encoders, it combines them using a simple linear strategy to obtain final node representations for fine-tuning. This approach is less flexible and effective than our proposed node- and dimension-adaptive fine-tuning strategy, as demonstrated in Theorem 1.
Appendix C Limitations
@@
While APF demonstrates strong performance across a wide range of benchmarks, it exhibits limitations in certain scenarios. Specifically, APF underperforms tree-based models such as XGBGraph on datasets like Amazon and Elliptic, where node features are highly heterogeneous and dominate over structural signals. This suggests that in such cases, tree-based models, known for their robustness to feature heterogeneity, may outperform deep learning-based GNNs [22]. Future work could explore hybrid architectures that better integrate rich tabular features with graph topology.
@@
Appendix D Proof of Theorem 1
@@
The proof is derived from [2], which analyzes the linear separability of a single graph convolution under a single CSBM model with only one pattern $CSBM(n,\boldsymbol{\mu},\boldsymbol{\nu},(p,q))$. We extend the analysis to our $ASBM(n_{a},n_{n},\boldsymbol{\mu},\boldsymbol{\nu},(p_{1},q_{1}),(p_{2},q_{2}),P_{a},P_{n})$.
Following [2], our analysis is based on the following assumptions: 1) The graph size $n$ is relatively large, with $\omega(d\log d)\leq n\leq\mathcal{O}(poly(d))$. 2) The graph is not too sparse, with $p_{1},q_{1},p_{2},q_{2}=\omega(\log^{2}(n)/n)$. The first assumption ensures that the number of nodes is at least quasilinearly large and at most polynomially large relative to the feature dimension. The second assumption ensures that the CSBM is not excessively sparse, maintaining a significant difference between the number of intra-class and inter-class edges. Additionally, we assume that the sizes of the homophilic and heterophilic node sets are approximately equal for theoretical tractability.
For simplicity, we use the random walk normalized adjacency matrix, $\boldsymbol{D}^{-1}\boldsymbol{A}$, as the low-pass filter, and its negative, $-\boldsymbol{D}^{-1}\boldsymbol{A}$, as the high-pass filter [2, 47, 45, 48, 23, 64]. Due to the Gaussian distribution of node features, the filtered node features also follow a Gaussian distribution. Specifically, for the low-pass filtered node features $\boldsymbol{\hat{X}}=\boldsymbol{D}^{-1}\boldsymbol{A}\boldsymbol{X}$, the means of nodes in different classes and patterns are given by [2]:
$\mathbb{E}(\boldsymbol{\hat{X}}_{i})=\left\{\begin{aligned} \frac{p_{1}\boldsymbol{\mu}+q_{1}\boldsymbol{\nu}}{p_{1}+q_{1}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{o},\\ \frac{q_{1}\boldsymbol{\mu}+p_{1}\boldsymbol{\nu}}{p_{1}+q_{1}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{o},\\ \frac{p_{2}\boldsymbol{\mu}+q_{2}\boldsymbol{\nu}}{p_{2}+q_{2}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{e},\\ \frac{q_{2}\boldsymbol{\mu}+p_{2}\boldsymbol{\nu}}{p_{2}+q_{2}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{e},\\ \end{aligned}\right.$ (15)
where $\mathcal{C}_{a}$ and $\mathcal{C}_{n}$ represent the abnormal and normal node sets, respectively, while $\mathcal{H}_{o}$ and $\mathcal{H}_{e}$ denote the homophilic and heterophilic node sets. The covariance matrix for $\boldsymbol{\hat{X}}_{i}$ is given by $\mathbb{C}(\boldsymbol{\hat{X}}_{i})=\frac{1}{d\boldsymbol{D}_{ii}}$. Based on Lemma 2 in [2], for any unit vector $\boldsymbol{w}$, we have
$\left|\left(\boldsymbol{\hat{X}}_{i}-\mathbb{E}(\boldsymbol{\hat{X}}_{i})\right)\cdot\mathbf{w}\right|=\mathcal{O}\left(\sqrt{\frac{\log n}{dn(p_{1}+q_{1})}}\right).$ (16)
When considering only the nodes in the homophily set $\mathcal{H}_{o}$, Lemma 3 in [2] shows that an optimal linear classifier can be found with the following parameters:
$\boldsymbol{w}^{*}=R\frac{\boldsymbol{\nu}-\boldsymbol{\mu}}{\|\boldsymbol{\mu}-\boldsymbol{\nu}\|},\quad b^{*}=-\frac{\langle\boldsymbol{\mu}+\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle}{2},$ (17)
where $R$ is the norm constraint, i.e., $\|\boldsymbol{w}\|\leq R$, and the distance between $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ is sufficiently large, with $\|\boldsymbol{\mu}-\boldsymbol{\nu}\|=\Omega(\frac{\log n}{\sqrt{dn(p_{1}+q_{1})}})$. This linear classifier can correctly separate homophily nodes $\mathcal{H}_{o}$. Specifically, for $i\in\mathcal{C}_{a}\cap\mathcal{H}_{o}$, we have:
$\langle\boldsymbol{\hat{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*}$ (18)
@@
$=\frac{\langle p_{1}\boldsymbol{\mu}+q_{1}\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle}{p_{1}+q_{1}}(1+o(1))+\mathcal{O}\left(\|\boldsymbol{w}^{*}\|\sqrt{\frac{\log n}{dn(p1+q1)}}\right)-\frac{\langle\boldsymbol{\mu}+\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle}{2}$
$=\frac{\langle 2p_{1}\boldsymbol{\mu}+2q_{1}\boldsymbol{\nu}-(p_{1}+q_{1})(\boldsymbol{\mu}+\boldsymbol{\nu}),\boldsymbol{w}^{*}\rangle}{2(p_{1}+q_{1})}(1+o(1))+o(\|\boldsymbol{w}^{*}\|)$
$=\frac{p_{1}-q_{1}}{2(p_{1}+q_{1})}\langle\boldsymbol{\mu}-\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle(1+o(1))+o(\|\boldsymbol{w}^{*}\|)$
$=-\frac{R(p_{1}-q_{1})}{2(p_{1}+q_{1})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))<0.$
Similarity, for $i\in\mathcal{C}_{n}\cap\mathcal{H}_{o}$, we obtain:
$\langle\boldsymbol{\hat{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*}=-\frac{R(q_{1}-p_{1})}{2(p_{1}+q_{1})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))>0.$ (19)
However, if we apply this linear classifier to the heterophilic node set $\mathcal{H}_{e}$, where $p_{2}<q_{2}$, all nodes are misclassified:

### v3

@@ 5 Conclusion
This paper tackles label scarcity and homophily disparity in GAD by introducing APF. APF first leverages Rayleigh Quotient-guided subgraph sampling and dual spectral filters to capture both semantic and anomaly-sensitive signals without supervision. During fine-tuning, a node- and dimension-adaptive fusion mechanism, together with anomaly-aware regularization, enhances the model’s ability to distinguish abnormal nodes under homophily disparity. Both theoretical analysis and extensive experiments on 10 benchmark datasets validate the effectiveness of our approach.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62306137).
Ethics Statement
We adhere to the ICLR Code of Ethics and have taken every measure to ensure that our research complies with the ethical guidelines set forth. Our work does not involve human subjects, nor does it raise any ethical concerns related to privacy or data usage. All datasets used in our experiments are publicly available, and we ensure that their release and use adhere to proper data-sharing policies. We have carefully selected the datasets and evaluation metrics to ensure fairness and transparency in our findings. No potential conflicts of interest, sponsorship biases, or ethical violations have been identified in our study. We commit to maintaining the highest standards of research integrity, and we are open to addressing any ethical concerns that may arise during the review process.
Reproducibility Statement
We have made efforts to ensure that this work is fully reproducible. All experiments were conducted with well-defined and publicly available datasets, and the models were implemented using standard frameworks. To facilitate the reproducibility of our results, we provide a detailed description of our experimental setup in the main text and Appendix, including hyperparameters, training protocols, evaluation methods, and the open-source link. We believe that all materials necessary for reproducing our experiments are clearly outlined.
Appendix A Use of Large Language Models
In accordance with the ICLR 2026 Policies on Large Language Model Usage, we disclose that Large Language Models (LLMs) were employed during the preparation of this manuscript. Specifically, LLMs were used to (i) check grammar and spelling, (ii) improve clarity and conciseness of sentences, (iii) format tables for consistency, and (iv) shorten certain passages to reduce redundancy. Importantly, LLMs were not used to generate research ideas, design experiments, analyze results, or draft new scientific content. The authors remain fully responsible for all claims, findings, and conclusions presented in this work.
Appendix B Detailed Preliminaries
@@
In a general scenario, we are given an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\boldsymbol{X})$, where $\mathcal{V}=\{v_{1},\cdots,v_{n}\}$ is the set of $n$ nodes, $\mathcal{E}=\{e_{ij}\}$ is the set of edges, and $e_{ij}=(v_{i},v_{j})$ represents an edge between nodes $v_{i}$ and $v_{j}$. We define $\boldsymbol{A}$ as the corresponding adjacency matrix and $\boldsymbol{D}$ as the diagonal degree matrix with $\boldsymbol{D}_{ii}=d_{i}=\sum_{j}\boldsymbol{A}_{ij}$. The neighbor set $\mathcal{N}_{i}$ of each node $v_{i}$ is given by $\mathcal{N}_{i}=\{v_{j}:e_{ij}\in\mathcal{E}\}$. For each node $v_{i}$, it has a $d$-dimensional feature vector $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$, and collectively the features of all nodes are denoted as $\boldsymbol{X}=(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n})^{\top}\in\mathbb{R}^{n\times d}$.
@@
GAD is formulated as a binary classification problem where anomalies are regarded as positive with label 1, while normal nodes are negative with label 0. Given $\mathcal{V}^{L}=\mathcal{V}_{a}\cup\mathcal{V}_{n}$, where $\mathcal{V}_{a}$ consists of labeled abnormal nodes and $\mathcal{V}_{n}$ comprises labeled normal nodes, along with their corresponding labels $\boldsymbol{y}^{L}$, the goal is to identify the anomalous status $\boldsymbol{\hat{y}}^{U}$ for the unlabeled nodes $\mathcal{V}^{U}=\mathcal{V}\setminus\mathcal{V}^{L}$. Usually, obtaining authentic labels is often costly, we assume that label information is only available for a small subset of nodes (i.e., $|\mathcal{V}^{L}|\ll|\mathcal{V}|$). Furthermore, there are usually significantly fewer abnormal nodes than normal nodes (i.e., $|\mathcal{V}_{a}|\ll|\mathcal{V}_{n}|$).
@@
Given the adjacency matrix $\boldsymbol{A}$, the graph Laplacian matrix is defined as $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{A}$, which is symmetric and positive semi-definite. Their symmetric normalized versions are noted as $\tilde{\boldsymbol{A}}=\boldsymbol{D}^{-1/2}\boldsymbol{A}\boldsymbol{D}^{-1/2}$ and $\tilde{\boldsymbol{L}}=\boldsymbol{I}-\boldsymbol{D}^{-1/2}\boldsymbol{A}\boldsymbol{D}^{-1/2}$. Its eigendecomposition is given by $\tilde{\boldsymbol{L}}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}$, where the columns of $\boldsymbol{U}\in\mathbb{R}^{n\times n}$ are orthonormal eigenvectors (graph Fourier basis), and $\boldsymbol{\Lambda}=\operatorname{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ contains the eigenvalues (frequencies), arranged such that $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$. Given the graph features $\boldsymbol{X}$ and a filter function $g(\cdot)$, the corresponding filtered features is thus defined as $\boldsymbol{\hat{X}}=\boldsymbol{U}g(\boldsymbol{\Lambda})\boldsymbol{U}^{\top}\boldsymbol{X}$.
@@
Typically, $\tilde{\boldsymbol{A}}$ acts as a low-pass filter with $g(\lambda)=1-\lambda$, while $\tilde{\boldsymbol{A}}$ and $\tilde{\boldsymbol{L}}$ serve as high-pass filters, with $g(\lambda)=\lambda-1$ and $g(\lambda)=\lambda$, respectively. In practice, a self-loop is often added to each node in the graph (i.e., $\boldsymbol{A}=\boldsymbol{A}+\boldsymbol{I}$) to alleviate numerical instabilities and improve performance Kipf and Welling (2017).
Appendix C Related Works
C.1 Graph Anomaly Detection
@@
GNN-based approaches have emerged as a promising paradigm for GAD, due to their strong ability to capture both complex structural and node attribute patterns in graph data. A comprehensive and up-to-date survey of deep GAD methods is provided in (Qiao et al., 2025b), while BOND (Liu et al., 2022) and GADBench (Tang et al., 2023) establish performance benchmarks for unsupervised and semi-/supervised GAD approaches, respectively.
@@
Unsupervised GAD approaches do not rely on labeled data for training and instead use unsupervised learning techniques to identify anomaly patterns in graph data. For instance, DOMINANT (Ding et al., 2019) employs a graph autoencoder to reconstruct both attributes and structure using GNNs. CoLA (Liu et al., 2021b) explores the consistency between anomalies and their neighbors across different contrastive views to assess node irregularity. VGOD (Huang et al., 2023) combines variance-based and attribute reconstruction models to detect anomalies in a balanced manner. TAM (Qiao and Pang, 2023) introduces local affinity as an anomaly measure, aiming to learn tailored node representations for GAD by maximizing the local affinity between nodes and their neighbors. GAD-EBM (Roy et al., 2023) evaluates the likelihood of normal and anomalous nodes to address GAD with an energy-based model. DiffGAD (Li et al., 2025) leverages diffusion sampling to infuse the latent space with discriminative content and introduces a content-preservation mechanism that retains valuable information across different scales for GAD.
@@
Semi-/supervised GAD approaches assume that labels for some normal and abnormal nodes are available for training. They aim to assign labels by learning a decision boundary between normal and abnormal nodes. For example, BWGNN (Tang et al., 2022) uses a Beta graph wavelet to learn band-pass filters that capture anomaly signals, while DSGAD (Zheng et al., 2025a) extends BWGNN with dynamic wavelets and feature fusion. AMNet (Chai et al., 2022) employs a restricted Bernstein polynomial parameterization to approximate filters in multi-frequency groups. CARE-GNN (Dou et al., 2020), PCGNN (Liu et al., 2021a), and GHRN (Gao et al., 2023) adaptively prune inter-class edges based on neighbor distributions or the graph spectrum. PMP (Zhuo et al., 2024) introduces a partitioned message-passing mechanism to handle homophilic and heterophilic neighbors independently. To address settings with limited labeled data, CGenGA (Ma et al., 2024) proposes a diffusion-based graph generation method to synthesize additional training nodes, while ConsisGAD (Chen et al., 2024b) incorporates learnable data augmentation to utilize the abundance of unlabeled data for consistency training. GGAD (Qiao et al., 2024) introduces a novel semi-supervised framework using only labeled normal nodes. SpaceGNN Dong et al. (2025) combines space projection, distance-aware propagation, and ensemble mechanisms across multiple latent spaces to improve generalization. It is worth noting that some approaches can also address label scarcity by leveraging auxiliary data, such as cross-network meta-learning (Meta-GDN (Ding et al., 2021b)) or cross-domain adaptation (Commander (Ding et al., 2021a), ACT (Wang et al., 2023)). In contrast, our proposed APF targets the single-graph setting where no external auxiliary networks or source domains are available, relying instead on mining intrinsic anomaly signals from unlabeled nodes within the target graph.
@@
Additionally, several works have explored GAD within the pre-training and fine-tuning paradigm. DCI (Wang et al., 2021) decouples representation learning and classification through a cluster-enhanced self-supervised learning task. Cheng et al. (2024) evaluates the performance of DGI (Veličković et al., 2019) and GraphMAE (Hou et al., 2022) for GAD, demonstrating the potential of leveraging graph pre-training to enhance GAD with limited supervision. However, most existing methods pre-train a uniform global low-pass filter (e.g., GCN (Kipf and Welling, 2017)) and then fine-tune a classifier on frozen node representations. The homophily disparity in GAD presents a significant challenge for directly applying these methods. To address this, we propose a pre-training and fine-tuning framework tailored for GAD.
In addition to the above work, PReNet (Pang et al., 2023) and NSReg (Wang et al., 2025) aim to identify anomalies whose patterns differ from labeled examples by enforcing strong normality modeling. GDN-AugAN (Zhou et al., 2023) enhances cross-dataset robustness through augmentation-based domain generalization. UniGAD (Lin et al., 2024), ARC (Liu et al., 2024a), UNPrompt (Niu et al., 2025), and AnomalyGFM (Qiao et al., 2025a) have explored foundation models for generalist GAD. While UniGAD (Lin et al., 2024) also employs a Rayleigh Quotient-based sampler, it uses the sampler as a unification tool to transform node/edge tasks into graph tasks for multi-level detection. In contrast, APF utilizes the sampler specifically to inject anomaly awareness during pre-training to address label scarcity in node anomaly detection.
C.2 Graph Pre-Training
@@
Graph pre-training has emerged as a promising paradigm for label-efficient learning (Cheng et al., 2023; Ju et al., 2024). These methods first learn universal knowledge from unlabeled data using self-supervised objectives, which are then transferred to tackle specific downstream tasks. Existing pre-training approaches can be broadly categorized into two groups: contrastive and non-contrastive approaches.
@@
Contrastive approaches typically follow the principle of mutual information maximization (Hjelm et al., 2019), where the objective functions contrast positive pairs against negative ones. For instance, DGI (Veličković et al., 2019) and DCI (Wang et al., 2021) focus on representation learning by maximizing the mutual information between node-level representations and a global summary representation. GRACE (Zhu et al., 2020), GCA (Zhu et al., 2021), and NS4GC (Liu et al., 2024b) learn node representations by pulling together the representations of the same node (positive pairs) across two augmented views, while pushing apart the representations of different nodes (negative pairs) across both views.
@@
Non-contrastive approaches, on the other hand, eliminate the need for negative samples. For example, CCA-SSG (Zhang et al., 2021) and G-BT (Bielak et al., 2022) aim to learn augmentation-invariant information while introducing feature decorrelation to capture orthogonal features. BGRL (Thakoor et al., 2022) and BLNN (Liu et al., 2024d) employ asymmetric architectures that learn node representations by predicting alternative augmentations of the input graph and maximizing the similarity between the predictions and their corresponding targets. GraphMAE Hou et al. (2022) focuses on feature reconstruction using a masking strategy and scaled cosine error. Additionally, SSGE (Liu et al., 2024c) minimizes the distance between the distribution of learned representations and the isotropic Gaussian distribution to promote the uniformity of node representations.
@@
However, the methods discussed above rely on low-pass GNN encoders that inherently smooth neighbor representations, leading to unsatisfactory performance on heterophilic abnormal nodes. Although a recent work, PolyGCL (Chen et al., 2024a), employs both low- and high-pass encoders, it combines them using a simple linear strategy to obtain final node representations for fine-tuning. This approach is less flexible and effective than our proposed node- and dimension-adaptive fine-tuning strategy, as demonstrated in Theorem 1.
Appendix D Limitations
@@
While APF demonstrates strong performance across a wide range of benchmarks, it exhibits limitations in certain scenarios. Specifically, APF underperforms tree-based models such as XGBGraph on datasets like Amazon and Elliptic, where node features are highly heterogeneous and dominate over structural signals. This suggests that in such cases, tree-based models, known for their robustness to feature heterogeneity, may outperform deep learning-based GNNs Grinsztajn et al. (2022). Future work could explore hybrid architectures that better integrate rich tabular features with graph topology.
@@
Appendix E Proof of Theorem 1
@@
We extend the argument of Baranwal et al. (2021) from a single-pattern $CSBM$ to the mixed-pattern, degree-corrected, and class-imbalanced $ASBM$ in Definition 1. Throughout the proof, we make the following standard assumptions:
  1. 1.
    the graph size is relatively large with $\omega(d\log d)\leq n\leq\mathcal{O}(poly(d))$;
  2. 2.
    sparsity level obeys $p_{1},q_{1},p_{2},q_{2}=\omega(\log^{2}n/n)$;
  3. 3.
    degree parameters are bounded and centered as $\theta_{\min}\leq\theta_{i}\leq\theta_{\max}$ with $\chi:=\theta_{\max}/\theta_{\min}=\mathcal{O}(1)$ and $\mathbb{E}[\theta_{i}\mid z_{i}]=1$ for $z_{i}\in\{a,n\}$;
  4. 4.
    class prior is imbalanced with $\pi_{a}=n_{a}/n\ll\pi_{n}=1-\pi_{a}$.
Following Baranwal et al. (2021); Ma et al. (2022); Mao et al. (2023); Han et al. (2024), we adopt the random-walk operator $\boldsymbol{S}=\boldsymbol{D}^{-1}\boldsymbol{A}$ as a low-pass filter and its negative $\boldsymbol{S}$ as a high-pass filter. Concretely, each node $v_{i}$ is assigned a spectral filter function $g(\cdot)$ depending on its pattern: $g(\lambda)=\lambda$ if $v_{i}\in\mathcal{H}_{o}$ (homophilic), and $g(\lambda)=-\lambda$ if $v_{i}\in\mathcal{H}_{e}$ (heterophilic). This yields the node-adaptive filtered features $\hat{\boldsymbol{X}}=\boldsymbol{U}g(\boldsymbol{\Lambda})\boldsymbol{U}^{\top}\boldsymbol{X}$.
Concentration under degree correction.
Let degree $d_{i}=\sum_{j}A_{ij}$ and recall that, conditional on $(z_{i},h_{i})$, we have $\mathbb{E}[A_{ij}\mid z_{i},z_{j},h_{i}]=\theta_{i}\theta_{j}\mathbf{B}^{(h_{i})}_{z_{i},z_{j}}$. Standard Bernstein (Chung and Lu, 2006; Lei and Rinaldo, 2015) bounds with bounded $\theta$ and $\omega(\log^{2}n/n)$ sparsity yield
$d_{i}=\theta_{i}\,n\big(\pi_{a}\beta_{z_{i},a}^{(h_{i})}+\pi_{n}\beta_{z_{i},n}^{(h_{i})}\big)\,(1\pm o(1)),$
where $\beta_{u,v}^{(h)}:=\mathbf{B}^{(h)}_{u,v}$. Hence for any node $v_{i}$ and any unit vector $\boldsymbol{w}$,
@@
$\hat{\boldsymbol{X}}_{i}=\pm\,\frac{1}{d_{i}}\sum_{j}A_{ij}\,\boldsymbol{X}_{j}=\pm\Big(\mathbb{E}[\hat{\boldsymbol{X}}_{i}\mid z_{i},h_{i}]+\boldsymbol{\xi}_{i}\Big),\quad\big|\langle\boldsymbol{\xi}_{i},\boldsymbol{w}\rangle\big|=\mathcal{O}\!\left(\sqrt{\frac{\log n}{d\,d_{i}}}\right)=\mathcal{O}\!\left(\sqrt{\frac{\log n}{d\,n\,\theta_{i}\,\kappa_{i}}}\right),$
with $\kappa_{i}:=\pi_{a}\beta_{z_{i},a}^{(h_{i})}+\pi_{n}\beta_{z_{i},n}^{(h_{i})}$. The sign ± corresponds to low- vs. high-pass filtering. Let the effective average connectivity under imbalance be
$\kappa_{\mathrm{eff}}:=\min\big\{\pi_{a}p_{1}+\pi_{n}q_{1},\ \pi_{a}q_{1}+\pi_{n}p_{1},\ \pi_{a}p_{2}+\pi_{n}q_{2},\ \pi_{a}q_{2}+\pi_{n}p_{2}\big\}.$
Using $\theta_{i}\geq\theta_{\min}$ and $\kappa_{i}\geq\kappa_{\mathrm{eff}}$, we obtain the uniform deviation bound (Baranwal et al., 2021)